Solve and graph on a number line.
|3x-8|>7
1
Expert's answer
2020-08-02T18:04:57-0400

Solution

To solve "|3x-8|>7, let\, 3x-8=7"

Add 8 to both sides as follows

"3x-8+8=7+8"

"3x=15"

divide by 3 both sides

"\frac{3x} 3=\frac {15} 3"

"x=5"

Again we let "3x-8=-7"

"3x-8+8=-7+8"

"3x=1"

"\frac{3x}3= \frac 13"

"x=\frac 1 3"

since "|3x-8|>7" is an absolute inequality, so now we have two solutions for x

"x=5" and

"x=\frac13"

Now the solution on a number line

"x<\frac 1 3"

and

"x>5"